Equivalents

Washington State Fisheries, Inc., processes salmon for various distributors. Two departments, processing and packaging, are involved. Data relating to tons of salmon sent to the processing department during May follow:

 




Percent Completed


Tons of Salmon


Direct Materials


Conversion

Work-in-process inventory, May 1

1,000


80

%


70

%

Work-in-process inventory, May 31

2,000


50

%


30

%

Started processing during May

7,000








 

Required:

1.    Calculate the number of tons completed and transferred out during the month.


2.    Calculate the number of equivalent units for both direct materials and conversion for the month of May, assuming that the company uses the weighted-average method.


3. How would your answer in requirement 2 change if the percentage of completion in ending inventory were as follows: direct materials 30%, conversion 40%?
Answer #1

Required:

1.    Calculate the number of tons completed and transferred out during the month.

1000+7000-2000= 6,000

2.    Calculate the number of equivalent units for both direct materials and conversion for the month of May, assuming that the company uses the weighted-average method.

Direct Materials= 6,000+ 1000= 7000      **2000*50%

Conversion= 6000+ 600= 6600     **2000*30%

3. How would your answer in requirement 2 change if the percentage of completion in ending inventory were as follows: direct materials 30%, conversion 40%?

Direct Materials= 6000+ 600= 6600   **2000*30%

Conversion= 6000+800= 6800            **2000*40%